# Prompts to Proxies: Emulating Human Preferences via a Compact LLM Ensemble

### v1

@@ Abstract
Large language models (LLMs) have demonstrated promise in emulating human-like responses across a wide range of tasks. In this paper, we propose a novel alignment framework that treats LLMs as agent proxies for human survey respondents, affording a cost-effective and steerable solution to two pressing challenges in the social sciences: the rising cost of survey deployment and the growing demographic imbalance in survey response data. Drawing inspiration from the theory of revealed preference, we formulate alignment as a two-stage problem: constructing diverse agent personas called endowments that simulate plausible respondent profiles, and selecting a representative subset to approximate a ground-truth population based on observed data. To implement the paradigm, we introduce P2P, a system that steers LLM agents toward representative behavioral patterns using structured prompt engineering, entropy-based sampling, and regression-based selection. Unlike personalization-heavy approaches, our alignment approach is demographic-agnostic and relies only on aggregate survey results, offering better generalizability and parsimony. Beyond improving data efficiency in social science research, our framework offers a testbed for studying the operationalization of pluralistic alignment. We demonstrate the efficacy of our approach on real-world opinion survey datasets, showing that our aligned agent populations can reproduce aggregate response patterns with high fidelity and exhibit substantial response diversity, even without demographic conditioning.
Keywords Preference modeling  $\cdot$ Value alignment  $\cdot$ Evaluation
@@ 1 Introduction
Pollsters are facing mounting challenges in the contemporary landscape, as public willingness to participate in surveys declines across regions and over time. In the United States, telephone response rates have dropped from 36% in the late 1990s to under 6% today Kennedy (2019), while major federal surveys such as the Current Population Survey (CPS) have experienced a 20-point drop since 2015 U.S. Bureau of Labor Statistics (2025). Likewise, across Europe Jabkowski and Cichocki (2025), response rates for national surveys in 36 European countries have fallen 0.8% per year on average between 1999 and 2018. This erosion of participation undermines statistical power and representativeness, particularly because the decline is not uniform across demographic groups. For instance, recent election polling failures have been traced to differential nonresponse: Republicans and independents are increasingly less likely to answer than Democrats Clinton et al. (2022); Parshall (2024). In the social sciences, this is a classic case of sample selection bias, which, if left unchecked, compromises both predictive accuracy and external validity of the survey-based inference. At the same time, the exorbitant cost of conducting large-scale surveys, due to respondent recruitment, compensation, and extensive piloting, places constraints on experimentation and iteration Kennedy (2015). These challenges highlight a pressing need for scalable, flexible, and theoretically grounded alternatives for modeling human preferences.
The epistemic challenge of capturing human preferences, long studied in the social sciences, now finds a structural parallel in large language model (LLM) alignment, where researchers aim to steer model behavior in line with human values and beliefs. Standard approaches, such as supervised fine-tuning (SFT) Tan et al. (2024) and reinforcement learning from human feedback (RLHF) Bai et al. (2022); Christiano et al. (2017), train models to reflect normative ideals of helpfulness, harmlessness, or truthfulness. While effective for building general-purpose assistants, these techniques are inherently monolithic: they produce a single dominant behavioral profile, making them ill-suited for modeling pluralistic human populations. Scaling them to reflect multiple value systems requires retraining or fine-tuning for each group, rendering the approach costly, brittle, and inflexible. Consequently, these pipelines struggle with tasks that demand population-level diversity, such as simulating public opinion or generating group-sensitive predictions.
To address this rigidity, alignment researchers have increasingly turned to post-training conditioning techniques that guide model behavior at inference time. A prominent example is the use of structured prompts, variously termed endowments Horton (2023), personas Castricato et al. (2025), or Silicon Samples Argyle et al. (2023), that encode sociodemographic attributes or behavioral dispositions. Prompted this way, LLMs can emulate individual-level responses with considerable fidelity Bisbee et al. (2024), making them attractive for constructing synthetic populations to pilot surveys or test hypotheses. However, this approach remains limited in key respects. First, it depends heavily on human-curated prompts or ground-truth demographic profiles, which may be unavailable or incomplete. Second, it often produces synthetic responses that lack the natural diversity of real populations: LLMs exhibit central-category bias Wang et al. (2024) and generate insufficient inter-agent variation Bisbee et al. (2024). Thus, while endowments reduce the need for retraining, they increase reliance on demographic templates and fall short of capturing true representational diversity.
These limitations reveal a common assumption in alignment: that behavioral fidelity hinges on demographic accuracy. We challenge this view and propose a new theoretical framework grounded in preference reconstruction. Rather than aligning each synthetic agent to a ground-truth demographic profile, we argue that effective emulation requires only the recovery of the aggregate preference structure of the target population. Drawing on revealed preference theory in economics, we treat each agent as a proxy decison-maker whose responses expose latent beliefs or utilities. Crucially, we do not seek to recover the true basis of individual human preferences. Instead, it suffices to identify a functional basis in the latent preference space, a diverse set of proxy agents that span the opinion manifold, and determine appropriate aggregation weights over them. This abstraction is reflected in human societies: demographically distinct populations may exhibit similar collective attitudes, even if their constituent beliefs diverge. Our approach thus enables a more data- and alignment-efficient route to survey emulation, moving beyond identity-conditioned generation toward flexible distributional modeling.
This perspective also clarifies and concretizes the goals of pluralistic alignment Sorensen et al. (2024), which advocates representing a spectrum of human values rather than optimizing for a singular objective. While compelling in principle, the pluralistic alignment literature has remained largely agnostic about how to measure pluralism or evaluate pluralistic systems. Our framework addresses this gap by proposing concrete criteria for measuring variation and fidelity, along with a mechanism to steer agent distributions efficiently under resource constraints so that the aggregate distribution of answers across the agent ensemble matches that of a given target population. In doing so, we operationalize distributional pluralistic alignment through a lens grounded in economic theory and suited for practical survey applications.
@@
  • Theoretical foundation. We formalize preference alignment as a two-stage process: constructing a functional basis of proxy agents and determining aggregation weights, drawing on revealed preference theory to recover population-level preferences.
  • Modular alignment system. We implement this theory via a modular system, Prompts to Proxies (P2P), that combines active endowment generation (to promote response diversity) with regression-based aggregation (to estimate preferences without demographic alignment).
  • Pluralistic alignment in practice. We provide the first operational framework for pluralistic alignment, introducing metrics for diversity and fidelity and propose scalable strategies to balance representational coverage and resource efficiency.
  • Bridge to empirical social science. Our framework enables social scientists to simulate and test preference theories by controlling agent endowments and tracking their effects on response distributions. It also supports survey design and can mitigate nonresponse bias, offering a cost-effective complement to traditional data collection.
  • Testbed for alignment research. The modular design provides a controlled environment for evaluating prompting strategies, enabling alignment researchers to explore steerability and value-sensitive modeling at scale.
@@ 2 Modeling Preferences via Functional Bases
Preference modeling under conventional lenses
Human preferences are latent, complex, and often inconsistent, posing a persistent challenge for formal modeling. Economists and social scientists have approached this challenge through multiple modeling frameworks. One prominent approach models humans as rational agents who maximize a stable utility function. While elegant in theory and mathematically tractable, these models often struggle to explain empirical anomalies in observed behavior Kahneman and Tversky (1979); Thaler (1999); Kahneman and Tversky (2000). A complementary framework, revealed preference theory Samuelson (1938); Mas-Colell et al. (1995), shifts the focus from explicit utility functions to observable choices. It asks whether a pattern of choices can be rationalized as if generated by a utility-maximizing agent without assuming a specific functional form. This idea has proven especially valuable in empirical settings, offering a way to assess whether behavior is consistent with stable preferences directly from observed data.
We take inspiration from the revealed preference perspective—not in its full rationality axioms, but in its core idea: preferences, though hidden, can be inferred from behaviors, such as responses to survey questions. This principle also underlies modern alignment pipelines such as SFT and RLHF, where models are steered toward human-preferred outputs using behavioral signals in the form of labeled data or feedback on generated responses.
Our framework
To formalize our approach, we zoom out from individual preferences and return to the level of survey analysis. Surveys are not designed to recover each respondent’s utility function; rather, they aim to capture aggregate population preferences, often through summary statistics such as response distributions or marginal means. This shift in perspective from individual to population invites a functional view (Figure 1): we can represent the population preference as a linear combination of individual preference vectors in a latent preference space. Responses to survey questions can then be viewed as mappings from this latent space to an observable response space. Crucially, this mapping is generally many-to-one: different underlying preferences may yield identical observed responses. Yet, as the number and diversity of responses increase, the expressive capacity of the response space improves. In the limit, a sufficiently rich battery of responses enables an increasingly precise identification of latent preferences. This perspective motivates our framework: Instead of recovering the ground-truth preference of each individual, we can construct a functional basis, a diverse set of proxy agents whose preferences span the latent space. By observing their responses and estimating the appropriate aggregation weights, which can be positive or negative, we can reconstruct the aggregate preference structure of the population.
@@
Alignment as a two-stage problem
Based on our theoretical framework, we can cast alignment as a two-stage problem. The first stage involves constructing a broad and expressive candidate basis of proxy agents, each representing a distinct vector in the latent preference space. For the basis to be effective, it must capture meaningful variation across the population. This objective aligns closely with recent calls for pluralistic alignment Sorensen et al. (2024), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal. However, we do not have direct access to individual preferences, and the prompt space is far too vast to explore in an unstructured manner. To address this, we adopt a structured prompting strategy based on attributes—interpretable factors whose variation is likely to influence preference expression. These attributes may be informed from theory (e.g., ideology, values, levels of altruism) or content (e.g., topical relevance). By combining and varying these attributes systematically, we generate a set of endowment-conditioned agents that span different positions in the preference space.
The second stage concentrates on designing an aggregation mechanism to reconstruct the population preference from proxy agents. We frame this as a supervised learning problem: given aggregate responses for a subset of survey questions, we estimate a set of aggregate weights over the proxy agents based on their individual response patterns. These weights are then used to predict responses on held-out questions, enabling out-of-sample evaluation. This setup allows us to examine whether the weighted agentic ensemble indeed captures the target population preference, rather than overfitting to the training set. We adopt regression-based methods for this purpose, as they align naturally with our assumption that population preferences emerge as linear combinations of individual vectors in the latent space. We describe implementation details, including attribute construction, endowment generation, and constrained regression methods, in the following section.
@@ 3 Prompts to Proxies (P2P)
To implement our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2). The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained variable section algorithms form the backbone. They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate the results of observed population-level surveys.
@@
3.1 Survey Splitting and Evaluation Protocol
To support the alignment procedure, we partition the full set of survey questions, denoted $\mathcal{Q}$, into three disjoint subsets: $\mathcal{Q}_{\mathrm{train}}$, $\mathcal{Q}_{\mathrm{valid}}$, and $\mathcal{Q}_{\mathrm{test}}$, corresponding to the training, validation, and test sets, respectively. Let $N=|\mathcal{Q}|$ denote the total number of questions, and $N_{\mathrm{train}}$, $N_{\mathrm{valid}}$, and $N_{\mathrm{test}}$ the cardinalities of the three subsets. P2P allows users to set the split sets and ratio flexibly.
For the alignment system, $\mathcal{Q}_{\mathrm{train}}$ is used for active endowment generation. $\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in regression-based aggregation. We reserve $\mathcal{Q}_{\mathrm{test}}$ as an ultimate test of the predictive accuracy of the selected agent ensemble, ensuring that the reconstructed preference has a decent degree of generalizability. For regression analysis, categorical questions are binarized into question-option pairs.
@@ 3.2 Attributes as Control Handles
As a crucial part of the active endowment generation logic, we use attributes to shape model preferences. This approach, though counter-intuitive at first glance, is in fact a deliberate reverse engineering design choice. In humans, attributes such as ideology, dispositions, or values are better understood as reflections or downstream manifestations of latent preferences. However, language models do not possess preferences in the human sense. As next-token prediction machines, their behavior is shaped by statistical associations learned from human text corpora. This distinction allows us to invert the causal direction: rather than treating attributes as reflections of internal preference states, we tap into the pattern matching nature of LLMs and use attributes as control handles to steer model behavior.
@@
To ensure that attributes effectively steer model preferences, we design a preset attribute bank that draws inspiration from theories across diverse social science disciplines, such as economics, political science, management science, and psychology. The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., values and beliefs, cognitive biases, or political ideology). These modes are further clustered into higher-level templates, which reflect different design purposes or theoretical orientations.
We define three primary templates. The core template contains general demographic descriptors and serves as the default baseline. The thematic template includes topic-specific modes such as economics, politics, culture and history. The theoretical template groups attributes derived from canonical frameworks such as Maslow’s hierarchy of needs Maslow (1943) and the Big Five personality traits John and Srivastava (1999). A condensed summary of these templates, representative modes, and attributes is shown in Table 1. These templates and modes are not fixed: they serve as a structured starting point but are fully extensible. Users can define new templates, introduce domain-specific modes, or modify existing attribute sets to suit different experimental contexts. By sampling and varying attributes from this flexible bank, we create prompts that instantiate distinct proxy agents occupying different regions of the latent preference space.
Template Representative Modes and Attributes
Core core: gender, race, ethnicity, education, income level,
Thematic economics: occupation, price sensitivity, wealth/debt,
politics: policy preferences, civic engagement, ideological orientation,
behavioral_biases: status quo bias, loss aversion, overconfidence bias,
Theoretical maslow: safety needs, self-actualization, love and belonging needs,
big_five: conscientiousness, neuroticism, extraversion,
Table 1: Structured view of the attribute bank, with representative modes and attributes grouped by template. Users can extend the bank with extra templates, modes, or attributes.
@@
While the attribute bank provides structured coverage, we also enable dynamic attribute discovery. We introduce two additional freeform templates that are specific to the survey under analysis. The survey template derives attributes by prompting a designated LLM, called attribute learner, to identify relevant decision-making factors based on the overall survey topic covered by the training set of questions, $\mathcal{Q}_{\mathrm{train}}$. The question template goes further by asking the learner to derive attributes tailored to a specific survey question. In both cases, the inferred attributes are appended to the attribute bank only for the duration of the survey run. This hybrid strategy strikes a balance between theory-driven and data-driven endowment generations, and enables human-AI co-discovery of salient preference descriptors.
With the attribute bank in place, we build an endowment generator using a dedicated LLM termed endowment model that takes a set of attributes as input and produces diverse agent profiles by varying the instantiation of each selected attribute.
3.3 Active Endowment Generation
To construct a diverse preference basis without exhaustively enumerating attribute combinations, we create an active endowment generation procedure to operate under resource constraints. It proceeds in iterative steps. At each step, new endowments are sampled, instantiated as agents to elicit responses to survey questions, and evaluated based on their contribution to response diversity. The remainder of this section describes the key components of this procedure, including our variability scoring strategy, adaptive sampling steps, and entropy-based patching mechanisms.

### v2

@@ Abstract
Large language models are increasingly used as proxies for human subjects in social science research, yet external validity requires that synthetic agents faithfully reflect the preferences of target human populations. We introduce preference reconstruction theory, a framework that formalizes preference alignment as a representation learning problem: constructing a functional basis of proxy agents and recovering population preferences through weighted aggregation. We implement this via Prompts to Proxies (P2P), a modular two-stage system. Stage 1 uses structured prompting with entropy-based adaptive sampling to construct a diverse agent pool spanning the latent preference space. Stage 2 employs L1-regularized regression to select a compact agent ensemble whose aggregate response distributions align with observed data from the target population. P2P requires no finetuning and no access to sensitive demographic data, incurring only API inference costs. We validate the approach on 14 waves of the American Trends Panel, achieving an average test MSE of 0.014 across diverse topics at approximately 0.8 USD per survey. We additionally test it on the World Values Survey, demonstrating its potential to generalize across locales. When stress-tested against an SFT-aligned baseline, P2P achieves competitive performance using less than 3% of the training data.
Preference Modeling, Value Alignment, Evaluation
@@ 1 Introduction
Large language models (LLMs) hold considerable promise as proxies for human subjects in social science research. From agent-based modeling of community dynamics (Park et al., 2023; Zhao et al., 2024a), to simulating survey responses (Argyle et al., 2023; Park et al., 2024) and prototyping policy interventions (Hou et al., 2025), synthetic agents afford scalability, reproducibility, and cost advantages unmatched by traditional human samples (Horton, 2023; Manning et al., 2024). However, realizing this promise hinges on solving a fundamental validity challenge: the preferences expressed by synthetic agents must faithfully reflect those of the target human population, at least when restricted to the topic domain. Without such alignment, conclusions drawn from synthetic experiments lack external validity (Shadish et al., 2002), limiting the scientific value of LLM-based social simulation.
This alignment challenge departs from conventional AI safety framing centered on steering a model toward a single set of desirable values, but shares motivation with the emerging notion of pluralistic alignment (Chen et al., 2025; Feng et al., 2024; Sorensen et al., 2024b) [^1] See Appendix E for an extensive review on this topic.. The goal here is distributional fidelity—ensuring that an ensemble of synthetic agents reproduces the heterogeneity of preference patterns observed in a target population. In reality, this challenge is compounded by an invisible resource barrier: while well-resourced industrial labs can deploy supervised fine-tuning to train personalized models on large annotated corpora, small academic institutions and independent researchers often face limited compute, limited data, and ethical and practical hurdles that preclude access to sensitive personal information. In this regard, existing approaches offer unsatisfying trade-offs: off-the-shelf LLMs trained via supervised fine-tuning (SFT) (Tan et al., 2024) and reinforcement learning from human feedback (RLHF) (Bai et al., 2022; Christiano et al., 2017) are easy to deploy but tend to collapse toward central tendencies in their training data, obscuring the preference heterogeneity (Feng et al., 2024; Kirk et al., 2024; Lee et al., 2024; Slocum et al., 2025); demographic conditioning methods require sensitive personal data and labor-intensive profile construction, yet still fail to guarantee distributional fidelity (Argyle et al., 2023; Castricato et al., 2025; Park et al., 2024; Bisbee et al., 2024).
In this work, we address this gap by introducing preference reconstruction theory, a framework that formalizes population preference alignment as a representation learning problem, inspired by revealed preference theory in economics. The core insight is that population-level preferences can be reconstructed by appropriately weighting a functional basis of proxy agents. We focus specifically on social survey data as our alignment signal, as they are publicly available, transparently documented, and designed to represent well-defined populations, yet typically withholding sensitive demographic information that would enable individual-level profile matching [^2] We expect the theory to extend to proprietary data in the format of categorical responses.. We operationalize this theory via P2P, an inference-based alignment system with a modular two-stage design. Stage 1 uses prompt engineering and entropy-based adaptive sampling to construct a diverse pool of agents with different response patterns, aimed to span the latent preference space of a topic domain. Stage 2 employs L1-regularized regression to select a representative and compact ensemble of agents whose response distributions match those of the target human population. This modularity allows each stage to be adapted or extended independently, accommodating different sampling strategies or aggregation methods as needed. We validate the design through simulation studies and evaluate P2P on the American Trends Panel and World Values Survey, demonstrating robust performance across topics and locales at low cost.
@@
  • Problem formulation. We formalize population preference alignment as a problem of representation learning, and introduce preference reconstruction theory, which states that learnable population preference can be represented using a functional basis of proxy agents with appropriately determined aggregation weights.
  • Modular alignment system. We operationalize preference reconstruction theory via Prompts to Proxies (P2P), a two-stage system that combines entropy-based adaptive sampling for diverse agent generation with L1-regularized regression for compact ensemble selection, requiring no fine-tuning, no demographic data, and modest resources to run.
  • Empirical validation. We evaluate P2P across 14 ATP waves and the World Values Survey, showing 43% improvement over prompting baselines, cross-locale generalization with theoretically grounded error analysis, and competitive fidelity against SFT approaches at less than 3% of the training data.
@@ 2 Modeling Preferences via Functional Bases
Alignment as representation learning
We formalize human-AI preference alignment as a representation learning problem, inspired by the revealed preference theory (Mas-Colell et al., 1995; Samuelson, 1938). Let $\mathcal{P}$ denote the latent human preference space. An individual human preference is represented as a vector in this space, $p\in\mathcal{P}$, which gets revealed via a decision-making mechanism $h$ into observed responses, denoted by vector $r\in\mathcal{R}\subset\mathbb{R}^{n}$. Similarly, let $\widehat{\mathcal{P}}$ be the latent model preference space. [^3] We intentionally define the model preference space $\widehat{\mathcal{P}}$ as an abstract space to accommodate the fact that both training and prompting can lead to different model preferences. In this paper, we use model and LLM interchangeably, as we contextualize alignment in the NLP context. A model preference is represented as $\hat{p}\in\widehat{\mathcal{P}}$, which manifests through the inference mechanism $f$ into the same observed response space $\mathcal{R}$. The representation learning process can thus be abstracted as:
[…]
where $f^{-1}$ is implemented via a learning algorithm [^4] Finding $\hat{p}$ such that $f(\hat{p})\approx r$, not a literal inverse.. We introduce learnability of preference in a revealed sense.
Definition 2.1.
Individual human preference is learnable in the revealed sense if $)h(\mathcal{P})\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)).
The definition using convex hull accommodates cases where individual human preference is not learnable by a single model but can be learned using an ensemble.
Learning population preference via survey data
While the above describes the learning process of an individual preference, alignment of large language models (LLMs) is usually done using responses from a human population. However, this target population is typically ill-defined for general-purpose LLMs, which have been through distinct stages of pretraining, supervised finetuning (SFT) and reinforcement learning from human feedback (RLHF), with data sources from different user groups. In this paper, we focus on preference alignment using aggregate responses in general social surveys, because these surveys are readily available, transparently documented, and designed to represent the views of well-defined target population. Formally, aggregate responses in a survey are defined as weighted averages of individual responses:
$r_{\mathrm{pop}}=\sum_{i=1}^{I}w_{i}h(p_{i})\ ,$ (2)
where $w_{i}$ is the weight of individuals with preference $p_{i}$ in the population, $I$. Accordingly, we extend our learnability definition to cover human population preference.
Definition 2.2.
Human population preference is learnable in the revealed sense if $)\mathrm{conv}\left(h(\mathcal{P})\right)\subset\mathrm{conv}(f(\mathcal{\widehat{P}}$)).
It follows readily that human population preference is learnable in the revealed sense if and only if individual human preference is learnable in the revealed sense (cf. Proposition A.4). Hence, in this paper we focus the discussion on population preference, with the expectation that the results extend to individual cases.
[…]
Observation 2.3.
If human population preference can be learned by a representative LLM agent, it can also be learned using a non-trivial [^6] A representative agent is by definition a degenerate ensemble. ensemble of LLM agents.
Observation 2.4.
If human population preference is learnable in the revealed sense, there could be cases where it can be learned by an ensemble but not by demographic conditioning nor a representative agent.
Observation 2.5.
If human population preference can learned by demographic conditioning, the ensemble that can learn it is not unique.
Preference reconstruction theory
Inspired by Observations 2.3-2.5, we propose preference reconstruction theory: instead of trying to find a one-to-one map between $p_{i}$ and $\hat{p}_{i}$, we can construct a functional basis—a set of proxy LLM agents with distinct preferences $\{\hat{p}_{j}\}_{j=1}^{J}$ spanning the relevant dimensions of human preferences—and then recover the human population preference through appropriate selection and weighting of these agents based on observational data. Figure 1 provides a functional perspective on preference reconstruction theory, together with a geometric interpretation of when the learnability assumption is violated (inadequate basis).
Alignment as a two-stage problem
Preference reconstruction via a functional basis can be formulated as a concrete learning problem:
$\min_{\{\hat{p}_{j},\tilde{w}_{j}\}_{j=1}^{J}}L(r_{\mathrm{pop}},\sum_{j=1}^{J}\tilde{w}_{j}f(x_{0};\hat{p}_{j}))\ ,$ (3)
where $L$ is a loss function.
Directly optimizing (3) is challenging, especially as the dimension $J$ is an optimizer per se.
As a solution strategy, we treat alignment as a two-stage problem with distinct stage goals. The first stage prioritizes the construction of a diverse set of proxy agents, measured in a revealed sense by their response variability. This goal aligns with the notion of pluralistic alignment (Sorensen et al., 2024b), which stresses the importance of representing a spectrum of human values rather than converging on a single ideal. The second stage concentrates on selecting a parsimonious ensemble from the set of candidate agents, based on the fit between ensemble aggregate responses and human ground truth. To avoid overfitting and test whether the ensemble has indeed learned the population preference, we partition survey questions into training and test sets, using the training set for selection and test set for validation. This solution strategy allows us to implement preference reconstruction modularly.
@@ 3 Prompts to Proxies (P2P)
To operationalize our preference reconstruction strategy in the survey context, we engineer a modular alignment system called Prompts to Proxies (P2P) (Figure 2). The system consists of two core components: (1) active endowment generation, powered by a dynamic attribute bank, and (2) regression-based aggregation, where constrained a variable selection algorithm forms the backbone. They concretize the two stages of alignment introduced in the previous section and enable the generation of a compact set of opinionated LLM agents to emulate population-level survey results.
3.1 Survey Splitting and Evaluation Protocol
To support the alignment procedure, the full set of survey questions denoted $\mathcal{Q}$ are partitioned into training, validation, and test subsets $\{\mathcal{Q}_{i}\}_{i=\mathrm{train},\mathrm{valid},\mathrm{test}}$. $\mathcal{Q}_{\mathrm{train}}$ is used in Stage 1 while $\mathcal{Q}_{\mathrm{train}}$ and $\mathcal{Q}_{\mathrm{valid}}$ are used in Stage 2. $\mathcal{Q}_{\mathrm{test}}$ serves as the validation to test preference generalizability.
@@ 3.2 Attributes as Control Handles
Language models, as next-token predictors, lack intrinsic preferences like humans. Their outputs are governed by statistical associations learned from text corpora and shaped by the user prompts. The prompt space is far too vast to explore in an unstructured manner. To address this, we adopt a structured prompting strategy based on attributes—interpretable factors, such as ideology, dispositions, or values, whose variation is likely to influence preference expression. P2P treats attributes as control handles for steering model behavior. [^7] This contrasts from the human case, where such attributes are typically viewed as reflections of latent preferences. P2P’s active endowment generation deliberately leverages this inversion: rather than inferring attributes from preferences, it uses attributes to shape the model’s expressed preferences.
@@
P2P constructs endowments (persona profiles) by drawing attributes from the attribute bank and instantiating each attribute with a specific value through a specialized LLM termed endowment model (cf. Appendix B.2.2) [^8] The use of endowment model ensures the created endowment remains coherent and interpretable to human readers.. The preset attribute bank for P2P lists example attributes taken from different social science disciplines, such as economics and political science. The bank is organized hierarchically: individual attributes are grouped into modes, each representing a coherent subject area (e.g., cognitive biases, political ideology). These modes are further clustered into higher-level templates, which reflect different design purposes or theoretical orientations. Figure 4 offers a structured view of the attribute bank and the endowment generation workflow.
@@
While the attribute bank ensures structured coverage, we also enable dynamic discovery through two freeform templates that are specific to the survey under analysis. The survey template prompts an LLM attribute learner to extract decision-making factors from the overall survey topic. The question template derives attributes for a specific survey question. These attributes are added to the bank only for the survey run. This hybrid approach balances theory- and data-driven methods, enabling human-AI co-discovery of key preference descriptors.
3.3 Stage 1: Active Endowment Generation
Stage 1 of P2P constructs a diverse preference basis without exhaustively enumerating attribute combinations. It begins by invoking an attribute learner to infer survey-specific attributes from the training questions, augmenting the attribute bank. Endowment generation then proceeds via adaptive sampling by mode: After initializing with equal endowments per mode, subsequent steps follow a bandit-style strategy—prioritizing modes that yield high response variability while exploring new modes for under-represented question items (see full Algorithm 1 in Appendix B.3).